OpenStudy (anonymous):

Solve the system of equations? 4x-y=8 and -8x+2y=-5 I think its no solution

OpenStudy (usukidoll):

solving a system of equations and "all real numbers" are unrelated. However, you can solve these problems by graphing, elimination, substitution, or using matrices.

OpenStudy (usukidoll):

I have to go for a bit... I"ll leave it up to the rest of the helpers on here.

OpenStudy (anonymous):

I tried using substitution and I got 0x=11 so...

OpenStudy (usukidoll):

we can use elimination method 4x-y=8 (first) -8x+2y=-5 (second) so let's multiply the entire first equation by 2 2(4x-y=8)

OpenStudy (anonymous):

8x-2y=16?

OpenStudy (usukidoll):

mhm 8x-2y=16 -8x+2y=-5 now what happens if we add?

OpenStudy (usukidoll):

what happens if we add these two equations together?

OpenStudy (anonymous):

The x's and y's get canceled out

OpenStudy (usukidoll):

yes and we're left with 0 = 11 is that a true statement or a false statement?

OpenStudy (anonymous):

false

OpenStudy (usukidoll):

yup because 0 is NEVER EQUAL to 11 when the x and y values cancel out, we have an inconsistent system. An inconsistent system means that there is no solution There is no x and no y value that will satisfy these equations. So we're done. :)

OpenStudy (anonymous):

Thank you